OpenStudy (anonymous):

how do you divide a number by a number when of them is a square root of the other.

OpenStudy (anonymous):

can you give an example?

OpenStudy (saifoo.khan):

E.g

OpenStudy (anonymous):

if you mean 25/sqrt(25) that's the same as 25^1/(25^1/2) so you subtract exponents and get sqrt(25) i believe

OpenStudy (anonymous):

which is 5

OpenStudy (anonymous):

Example \[1/\sqrt{2}\]?

OpenStudy (anonymous):

\[1/\sqrt{2}\]\[(1/\sqrt{2}) * (\sqrt{2}/\sqrt{2})\]\[\sqrt{2}/2\]

OpenStudy (anonymous):

yes. you answered completely right

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

it would be better if you had the example for us to see

OpenStudy (anonymous):

its fine. you answered fine. you worked out the problem and worked around the parts you were missing. this shows your true intelligence.